# Why High-rank Neural Networks Generalize?: An Algebraic Framework with RKHSs

### v1

@@ 1 Introduction
Understanding the generalization property of deep neural networks has been one of the biggest challenges in the machine learning community. It describes how the model can fit unseen data. Classically, the generalization error is bounded using the VC-dimension theory (Harvey et al., 2017; Anthony & Bartlett, 2009). Norm-based  (Neyshabur et al., 2015; Bartlett et al., 2017; Golowich et al., 2018; Neyshabur et al., 2018; Wei & Ma, 2019; Li et al., 2021; Ju et al., 2022; Weinan E et al., 2022) and compression-based (Arora et al., 2018; Suzuki et al., 2020) bounds have also been investigated. The norm-based bounds depend on the matrix $(p,q)$ norm of the weight matrices, and the compression-based bounds are derived by investigating how much the networks can be compressed. These bounds imply that low-rank weight matrices and weight matrices with small singular values, i.e., nearly low-rank matrices, have good effects for generalization.
@@
On the other hand, phenomena in which models with high-rank weight matrices generalize well have been empirically observed (Goldblum et al., 2020). Since the norm-based and compression-based bounds focus only on the low-rank and nearly low-rank cases, they cannot describe these phenomena. To theoretically describe these phenomena, the Koopman-based bound was proposed (Hashimoto et al., 2024). Koopman operators are linear operators that describe the compositions of functions, which are essential structures of neural networks. This bound is described by the ratio of the norm to the determinant of the weight matrix. Since the determinant factor appears in the denominator of the bound, even if the weight matrices are high rank and have large singular values, this bound can be small. The Koopman-based bound theoretically sheds light on why neural networks with high-rank weight matrices generalize well.
@@
However, the existing analysis for the Koopman-based bound strongly depends on the smoothness of models and the unboundedness of the data space, which excludes realistic models with bounded data space and with activation functions such as the hyperbolic tangent, sigmoid, and ReLU-type nonsmooth functions. In addition, the dependency of the bound on the activation function is not clear. The bound is described by the norms of the Koopman operators with respect to the activation functions, which is hard to evaluate in many cases.
@@
In this paper, we propose a new Koopman-based bound that resolves the issues of the existing Koopman-based bounds. Similar to the existing Koopman-based bounds, the proposed bound describes why high-rank neural networks generalize well but can be applied to a wider range of models than the existing Koopman-based bounds. We introduce a kernel function defined on the parameter space using linear operators on a Hilbert space to which models belong. This kernel function allows us to construct a reproducing kernel Hilbert space (RKHS) that describes realistic deep models with nonsmooth activation function and bounded data space. For example, we can set the Hilbert space as the $L^{2}$ function space on a bounded data space, which enables us to analyze models that the existing Koopman-based bounds cannot describe. We use the Rademacher complexity to derive generalization bounds. Using the reproducing property of the RKHS, we can bound the Rademacher complexity with the operator norms of the linear operators. For linear operators, we use group representations and Koopman operators. We first focus on algebraic representations of models using group representations. A typical example is the representation of the affine group, which describes invertible neural networks. We then focus on representations using Koopman operators with respect to the weight matrices, which describe neural networks with non-constant width.
@@
For $d\in\mathbb{N}$ and a Lebesgue measure space $\mathcal{X}\subseteq\mathbb{R}^{d}$, let $L^{2}(\mathcal{X})$ be the space of complex-valued squared Lebesgue-integrable functions on $\mathcal{X}$. We denote by $\mu_{\mathcal{X}}$ the Lebesgue measure on $\mathcal{X}$. For a Hilbert space $\mathcal{H}$, let $\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}}$ be the inner product in $\mathcal{H}$. We omit the subscript $\mathcal{H}$ when it is obvious. We denote by $B(\mathcal{H})$ the space of bounded linear operators on $\mathcal{H}$. All the technical proofs are documented in Appendix A.
@@ 2 Preliminaries
We review the fundamental notions required in this paper.
@@ 2.1 Koopman operator
The following lemma is regarding the boundedness of well-known elementwise activation functions, which is defined as $\sigma([x_{1},\ldots,x_{d}])=[\tilde{\sigma}(x_{1}),\cdots,\tilde{\sigma}(x_{d})]$ for a map $\tilde{\sigma}$$\mathbb{R}\to\mathbb{R}$.
@@ 3.1 Algebraic representation of deep models with group representations
Let $G=GL(d)\ltimes\mathbb{R}^{d}$ be the affine group and $\mathcal{H}=L^{2}(\mathbb{R}^{d})$. Let $\rho:G\to B(\mathcal{H})$ be the representation of $G$ on $\mathcal{H}$ defined as $\rho(g)h(x)=|\det W|^{1/2}h(W(x-b))$ for $g=(W,b)\in G$, $h\in L^{2}(\mathbb{R}^{d})$, and $x\in\mathbb{R}^{d}$. Note that $\rho$ is an irreducible unitary representation. In addition, let $v\in L^{2}(\mathbb{R}^{d})$ be the final nonlinear transformation, $\sigma_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be an activation function satisfying Assumption 1, and $A_{l}=K_{\sigma_{l}}$ be the Koopman operator with respect to $\sigma_{l}$ for $l=1,\ldots,L-1$. For example, $\sigma_{l}$ is the elementwise Leaky ReLU. Then, the deep model (1) is
@@
where $p_{c,x}\in\mathcal{H}$ with $\|p_{c,x}\|=1$ for $c>0$ and $x\in\mathcal{X}_{0}$. This regularization is required to technically derive the Rademacher complexity bound using the framework of RKHSs. However, as the following example indicates, the regularized model (2) sufficiently approximates the original model (1).
@@
Consider the same setting in Example 1. Let $p_{c,x}(y)=\sqrt{c/\pi}^{d}e^{-c\|y-x\|^{2}}$ for $c>0$ and $x\in\mathcal{X}_{0}$. Then, $p_{c,x}\in L^{2}(\mathbb{R}^{d})$ and $\|p_{c,x}\|=1$. We have
@@ 3.2 RKHS for analyzing deep models
We use the Rademacher complexity to derive a generalization bound. According to Theorem 3.5 in Mohri et al. (2018), the generalization bound is described by the Rademacher complexity. Thus, if we obtain a Rademacher complexity bound, then we can also bound the generalization error. To derive a Rademacher complexity bound, we apply the framework of RKHSs. The Hilbert space $\mathcal{H}$ to which the modes belong does not always have the reproducing property. Indeed, a typical example of $\mathcal{H}$ is $L^{2}(\mathbb{R}^{d})$ as we discussed in Example 1. Thus, we consider an RKHS that is a function space on the parameter space $G$ and isomorphic to a subspace of $\mathcal{H}$. We can regard the deep model on the data space $\mathcal{X}_{0}$ as a function on $G$ through this isomorphism and make use of the reproducing property on $G$. We define the following positive definite kernel $k:(G\times\cdots\times G)\times(G\times\cdots\times G)\to\mathbb{C}$ to construct an RKHS to analyze the deep model (1):
@@ 4 Rademacher Complexity Bound
We apply the isomorphism in Proposition 8 to derive a Rademacher complexity bound with the aid of the reproducing property in the RKHS $\mathcal{R}_{k}$. If $p_{c,x}\in\mathcal{K}$, Eq. (2) implies $F_{c}(\cdot,x)=\iota(p_{c,x})\in\mathcal{R}_{k}$ for $x\in\mathcal{X}_{0}$ and $c>0$. Thus, we can apply the reproducing property with respect to the model $F_{c}(\cdot,x)$. Let $\Omega$ be a probability space equipped with a probability measure $P$. Let $S\in\mathbb{N}$ be the sample size, $x_{1},\ldots,x_{S}\in\mathcal{X}_{0}$, and $\epsilon_{1},\ldots,\epsilon_{S}:\Omega\to\mathbb{C}$ be i.i.d. Rademacher variables. For a measurable function $\epsilon:\Omega\to\mathbb{C}$ , we denote by $\mathrm{E}[\epsilon]$ the integral $\int_{\Omega}\epsilon(\omega)\mathrm{d}P(\omega)$. The empirical Rademacher complexity $\hat{R}(\mathcal{F},x_{1},\ldots,x_{S})$ of a function class $\mathcal{F}$ is defined as $\hat{R}(\mathcal{F},x_{1},\ldots,x_{S})=\mathrm{E}[\sup_{F\in\mathcal{F}}\sum_{s=1}^{S}F(x_{s})\epsilon_{s}]/S$. We denote by $\mathcal{F}_{c}$ the function class $\{F_{c}(g_{1},\ldots,g_{L},\cdot)\,\mid\,g_{1},\ldots,g_{L}\in G\}$. The Rademacher complexity of $\mathcal{F}_{c}$ is upper bounded as follows.
@@
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})\leq\frac{\|A_{1}\|\cdots\|A_{L-1}\|\|v\|}{\sqrt{S}}.$
@@ 4.1 Invertible neural networks
$\hat{R}(\mathcal{NN}_{c},x_{1},\ldots,x_{S})\leq\frac{\|v\|\prod_{l=1}^{L-1}\|A_{l}\|}{\sqrt{S}}\sup_{|\det{W_{l}}|^{-1/2}\leq D}\prod_{l=1}^{L}|\det W_{l}|^{-1/2}.$
@@ 5.1 Algebraic representation of deep models with Koopman operators
where $p_{c,x}\in\mathcal{H}_{0}$ with $\|p_{c,x}\|=1$ for $x\in\mathcal{X}_{0}$ and $c>0$. We also define a positive definite kernel $k:(\Theta_{1}\times\cdots\times\Theta_{L})\times(\Theta_{1}\times\cdots\times\Theta_{L})\to\mathbb{C}$ to construct an RKHS to analyze the deep model (1):
@@ 5.2 Neural network with injective weight matrices
Let $d_{l}\in\mathbb{N}$ and $\Theta_{l}=\{W\in\mathbb{C}^{d_{l}\times d_{l-1}}\,\mid\,W\mbox{ is injective}\}$. Let $\mathcal{X}_{0}\subset\mathbb{R}^{d_{0}}$, $W_{l}\sigma_{l-1}(W_{l-1}\cdots W_{2}\sigma_{1}(W_{1}\mathcal{X}_{0}))\subseteq\tilde{\mathcal{X}_{l}}\subseteq\mathbb{R}^{d_{l}}$, and $\sigma_{l}(W_{l}\cdots W_{2}\sigma_{1}(W_{1}\mathcal{X}_{0}))\subseteq\mathcal{X}_{l}\subseteq\mathbb{R}^{d_{l}}$ that satisfy $\mu_{\mathbb{R}^{d_{l}}}(\mathcal{X}_{l})>0$ and $\mu_{\mathbb{R}^{d_{l}}}(\tilde{\mathcal{X}}_{l})>0$. Let $\tilde{\mathcal{H}}_{l}=L^{2}(\tilde{\mathcal{X}}_{l})$, $\mathcal{H}_{l}=L^{2}(\mathcal{X}_{l})$, and $\eta_{l}(W_{l})=K_{W_{l}}$ be the Koopman operator from $\tilde{\mathcal{H}}_{l}$ to $\mathcal{H}_{l-1}$ with respect to $W_{l}$. In addition, let $A_{l}=K_{\sigma_{l}}$ be the Koopman operator from $\mathcal{H}_{l}$ to $\tilde{\mathcal{H}}_{l}$ with respect to an activation function $\sigma_{l}:\tilde{\mathcal{X}}_{l}\to\mathcal{X}_{l}$ that satisfies Assumption 1. Then, we have
@@
Let $D>0$ and $\mathcal{F}_{c}=\{F_{c}(\theta_{1},\ldots,\theta_{L},\cdot)\,\mid\,|\det{W_{1}^{*}W_{1}}|^{-1/4},\ldots,|\det{W_{L}^{*}W_{L}}|^{-1/4}\leq D\}$. In the same way as in Theorem 11, we obtain the following bound.
@@
Assume $p_{c,x}\in\mathcal{K}$ for $x\in\mathcal{X}_{0}$. Let $f_{l}=v\circ W_{L}\circ\sigma_{L-1}\circ\cdots\circ W_{l+1}\circ\sigma_{l}$. Let
for $h\in\tilde{\mathcal{H}}_{l}$. Then, we have
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})\leq\sup_{|\det{W_{l}^{*}W_{l}}|^{-1/4}\leq D}\frac{\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\alpha(f_{l})}{\sqrt{S}\prod_{l=1}^{L}|\det W_{l}^{*}W_{l}|^{1/4}},$ (3)

### v2

@@ 1 Introduction
Understanding the generalization property of deep neural networks has been one of the biggest challenges in the machine learning community. The generalization property describes how the model can fit unseen data. Classically, the generalization error is bounded using the VC-dimension theory (Harvey et al., 2017; Anthony & Bartlett, 2009). Norm-based  (Neyshabur et al., 2015; Bartlett et al., 2017; Golowich et al., 2018; Neyshabur et al., 2018; Wei & Ma, 2019; Li et al., 2021; Ju et al., 2022; Weinan E et al., 2022) and compression-based (Arora et al., 2018; Suzuki et al., 2020) bounds have also been investigated. The norm-based bounds depend on the matrix $(p,q)$ norm of the weight matrices, and the compression-based bounds are derived by investigating how much the networks can be compressed. These bounds imply that low-rank weight matrices and weight matrices with small singular values, i.e., nearly low-rank matrices, have good effects for generalization. See Appendix C for more details about the existing bounds.
@@
On the other hand, phenomena in which models with weight matrices that are high-rank and have large singular values generalize well have been empirically observed (Goldblum et al., 2020). Since the norm-based and compression-based bounds focus only on the low-rank and nearly low-rank cases, they cannot describe these phenomena. To theoretically describe these phenomena, the Koopman-based bound was proposed (Hashimoto et al., 2024). Koopman operators are linear operators that describe the compositions of functions, which are essential structures of neural networks. This existing bound is described by the ratio of the norm to the determinant of each weight matrix as
${\color[rgb]{0,0,0}O\bigg(\prod_{l=1}^{L}\frac{G_{l}\|K_{\sigma_{l}}\|_{H_{l}}\|W_{l}\|^{s_{l-1}}}{\sqrt{S}\operatorname{det}(W_{l}^{*}W_{l})^{1/4}}\bigg),}$ (1)
where $S$ is the sample size, $s_{l}$ represents the smoothness of the $l$th layer, $G_{l}$ is a factor determined by the $l\sim L$th layers, $K_{\sigma_{l}}$ is the Koopman operator with respect to the activation function $\sigma_{l}$, and $\|\cdot\|_{H_{l}}$ represents the operator norm in a Sobolev space $H_{l}$. Since the determinant factor appears in the denominator of the bound, even if the weight matrices are high rank and have large singular values, this bound can be small. The Koopman-based bound theoretically sheds light on why neural networks with high-rank weight matrices generalize well.
@@
However, the existing analysis for the Koopman-based bound strongly depends on the smoothness of models and the unboundedness of the data space, which excludes realistic models with bounded data space and with activation functions such as the hyperbolic tangent, sigmoid, and ReLU-type nonsmooth functions. In addition, the dependency of the bound on the activation function is not clear. In fact, the factors $\|K_{\sigma_{l}}\|_{H_{l}}$ and $G_{l}$ in the bound (1) is hard to evaluate in many cases.
In this paper, we propose a new Koopman-based bound that resolves the issues of the existing Koopman-based bounds. The proposed bound is described as
${\color[rgb]{0,0,0}O\bigg(\prod_{l=1}^{L}\frac{G_{l}\|K_{\sigma_{l}}\|_{\mathcal{L}_{l}}}{\sqrt{S}\operatorname{det}(W_{l}^{*}W_{l})^{1/4}}\bigg),}$
where $\|\cdot\|_{\mathcal{L}_{l}}$ is the operator norm in a $L^{2}$ function space. Similar to the existing Koopman-based bounds, the proposed bound describes why high-rank neural networks generalize well. On the other hand, the difference of the function space $\mathcal{L}_{l}$ from $H_{l}$ gives a significant benefit to the proposed bound. We note that $\mathcal{L}_{l}$ is larger than $H_{l}$, and $\mathcal{L}_{l}$ enables us to analyze nonsmooth deep models and bounded data space. In addition, it enables us to evaluate the factors $\|K_{\sigma_{l}}\|_{\mathcal{L}_{l}}$ and $G_{l}$ easily (see Lemmas 24) and understand the effect of the activation functions on the deep model. As a result, the proposed bound significantly improves the existing bound in the sense that it can be applied to a wider range of models and enables us to understand the models well.
@@
To achieve the above improvement, we introduce a kernel function defined on the parameter space using linear operators on a Hilbert space to which models belong. This kernel function allows us to construct a reproducing kernel Hilbert space (RKHS) that describes realistic deep models with nonsmooth activation function and bounded data space. We use the Rademacher complexity to derive generalization bounds. The Rademacher complexity measures the complexity of the model, which also describe the generalization property. Using the reproducing property of the RKHS, we can bound the Rademacher complexity with the operator norms of the linear operators. For linear operators, we use group representations and Koopman operators. We first focus on algebraic representations of models using group representations. A typical example is the representation of the affine group, which describes invertible neural networks. We then focus on representations using Koopman operators with respect to the weight matrices, which describe neural networks with non-constant width. We schematically show the summary of the framework of the existing and proposed Koopman-based bounds in Figure 1.
@@
Figure 1: Summary of the framework of the existing and proposed Koopman-based bounds
@@
For $d\in\mathbb{N}$ and a Lebesgue measure space $\mathcal{X}\subseteq\mathbb{R}^{d}$, let $L^{2}(\mathcal{X})$ be the space of complex-valued squared Lebesgue-integrable functions on $\mathcal{X}$. We denote by $\mu_{\mathcal{X}}$ the Lebesgue measure on $\mathcal{X}$. For a Hilbert space $\mathcal{H}$, let $\left\langle\cdot,\cdot\right\rangle_{\mathcal{H}}$ be the inner product in $\mathcal{H}$. We omit the subscript $\mathcal{H}$ when it is obvious. We denote by $B(\mathcal{H}_{1},\mathcal{H}_{2})$ be the space of bounded linear operators from $\mathcal{H}_{1}$ to $\mathcal{H}_{2}$. In particular, we denote $B(\mathcal{H},\mathcal{H})=B(\mathcal{H})$. All the technical proofs are in Appendix A.
@@ 2.1 Koopman operator
The following lemma is regarding the boundedness of well-known elementwise activation functions defined as
$\sigma([x_{1},\ldots,x_{d}])=[\tilde{\sigma}(x_{1}),\cdots,\tilde{\sigma}(x_{d})]$
for a map $\tilde{\sigma}$$\mathbb{R}\to\mathbb{R}$.
@@ 3.1 Algebraic representation of deep models with group representations
Let $G=GL(d)\ltimes\mathbb{R}^{d}$ be the affine group and $\mathcal{H}=L^{2}(\mathbb{R}^{d})$. Here, $GL(d)$ is the group of $d$ by $d$ invertible matrices. Let $\rho:G\to B(\mathcal{H})$ be the representation of $G$ on $\mathcal{H}$ defined as $\rho(g)h(x)=|\det W|^{1/2}h(W(x-b))$ for $g=(W,b)\in G$, $h\in L^{2}(\mathbb{R}^{d})$, and $x\in\mathbb{R}^{d}$. Note that $\rho$ is an irreducible unitary representation. In addition, let $v\in L^{2}(\mathbb{R}^{d})$ be the final nonlinear transformation, $\sigma_{l}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be an activation function satisfying Assumption 1, and $A_{l}=K_{\sigma_{l}}$ be the Koopman operator with respect to $\sigma_{l}$ for $l=1,\ldots,L-1$. For example, $\sigma_{l}$ is the elementwise Leaky ReLU. Then, the deep model (2) is
@@
where $p_{c,x}\in\mathcal{H}$ for $c>0$ and $x\in\mathcal{X}_{0}$. We assume for any $c>0$, there exists $E(c)>0$ such that $\|p_{c,x}\|^{2}\leq E(c)$. This regularization is required to technically derive the Rademacher complexity bound using the framework of RKHSs. However, as the following example indicates, the regularized model (3) sufficiently approximates the original model (2).
@@
Consider the same setting in Example 1. Let $p_{c,x}(y)=(c/\pi)^{d/2}\mathrm{e}^{-c\|y-x\|^{2}}$ for $c>0$ and $x\in\mathcal{X}_{0}$. Then, $p_{c,x}\in L^{2}(\mathbb{R}^{d})$ and $\|p_{c,x}\|^{2}=(2c/\pi)^{d/2}$. Since $p_{c,x}$ goes to the Dirac delta function centered at $x$ as $c\to\infty$, we have
@@ 3.2 RKHS for analyzing deep models
We use the Rademacher complexity to derive a generalization bound. According to Theorem 3.5 in Mohri et al. (2018), the generalization error is bounded by the Rademacher complexity. Thus, if we obtain a Rademacher complexity bound, then we can also bound the generalization error. To derive a Rademacher complexity bound, we apply the framework of RKHSs. The Hilbert space $\mathcal{H}$ to which the models belong does not always have the reproducing property. Indeed, a typical example of $\mathcal{H}$ is $L^{2}(\mathbb{R}^{d})$ as we discussed in Example 1. Thus, we consider an RKHS that is a function space on the parameter space $G$ and isomorphic to a subspace of $\mathcal{H}$. We can regard the deep model on the data space $\mathcal{X}_{0}$ as a function on $G$ through this isomorphism and make use of the reproducing property on $G$. Here, the isomorphism ensures that the mathematical structure of the RKHS is the same as the subspace of $\mathcal{H}$. We define the following positive definite kernel $k:(G\times\cdots\times G)\times(G\times\cdots\times G)\to\mathbb{C}$ to construct an RKHS to analyze the deep model (2):
@@ 4 Rademacher Complexity Bound
We apply the isomorphism in Proposition 8 to derive a Rademacher complexity bound with the aid of the reproducing property in the RKHS $\mathcal{R}_{k}$. If $p_{c,x}\in\mathcal{K}$, Eq. (3) implies $F_{c}(\cdot,x)=\iota(p_{c,x})\in\mathcal{R}_{k}$ for $x\in\mathcal{X}_{0}$ and $c>0$. Thus, we can apply the reproducing property with respect to the model $F_{c}(\cdot,x)$. Let $\Omega$ be a probability space equipped with a probability measure $P$. Let $S\in\mathbb{N}$ be the sample size, $x_{1},\ldots,x_{S}\in\mathcal{X}_{0}$, and $\epsilon_{1},\ldots,\epsilon_{S}:\Omega\to\mathbb{C}$ be i.i.d. Rademacher variables (random variables following the uniform distribution on $\{-1,1\}$). For a measurable function $\epsilon:\Omega\to\mathbb{C}$, we denote by $\mathrm{E}[\epsilon]$ the integral $\int_{\Omega}\epsilon(\omega)\mathrm{d}P(\omega)$. The empirical Rademacher complexity $\hat{R}(\mathcal{F},x_{1},\ldots,x_{S})$ of a function class $\mathcal{F}$ is defined as $\hat{R}(\mathcal{F},x_{1},\ldots,x_{S})=\mathrm{E}[\sup_{F\in\mathcal{F}}\sum_{s=1}^{S}F(x_{s})\epsilon_{s}]/S$. We denote by $\mathcal{F}_{c}$ the function class $\{F_{c}(g_{1},\ldots,g_{L},\cdot)\,\mid\,g_{1},\ldots,g_{L}\in G\}$. The Rademacher complexity of $\mathcal{F}_{c}$ is upper bounded as follows.
@@
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})\leq\frac{\|A_{1}\|\cdots\|A_{L-1}\|\|v\|{\color[rgb]{0,0,0}E(c)}}{\sqrt{S}}.$
@@
Remark 13
If $p_{c,x}(y)=(c/\pi)^{d/2}\mathrm{e}^{-c\|y-x\|^{2}}$, then we have $E(c)=({2c}/{\pi})^{d/2}$. Combining with the discussion in Example 3, we can see that there is a tradeoff between $F_{c}$ being close to the original model $f$ and the constant $E(c)$ becoming large.
@@ 4.1 Invertible neural networks
$\hat{R}(\mathcal{NN}_{c},x_{1},\ldots,x_{S})\leq\frac{{\color[rgb]{0,0,0}E(c)}\|v\|\prod_{l=1}^{L-1}\|A_{l}\|}{\sqrt{S}}\sup_{|\det{W_{l}}|^{-1/2}\leq D}\prod_{l=1}^{L}|\det W_{l}|^{-1/2}.$
@@ 5.1 Algebraic representation of deep models with Koopman operators
where $p_{c,x}\in\mathcal{H}_{0}$ for $x\in\mathcal{X}_{0}$ and $c>0$ with $\|p_{c,x}\|\leq E(c)$ for $E(c)>0$. We also define a positive definite kernel $k:(\Theta_{1}\times\cdots\times\Theta_{L})\times(\Theta_{1}\times\cdots\times\Theta_{L})\to\mathbb{C}$ to construct an RKHS to analyze the deep model (2):
@@ 5.2 Neural network with injective weight matrices
Let $d_{l}\in\mathbb{N}$ and $\Theta_{l}=\{W\in\mathbb{C}^{d_{l}\times d_{l-1}}\,\mid\,W\mbox{ is injective}\}$. Let $\mathcal{X}_{0}\subset\mathbb{R}^{d_{0}}$, $W_{l}\sigma_{l-1}(W_{l-1}\cdots W_{2}\sigma_{1}(W_{1}\mathcal{X}_{0}))\subseteq\tilde{\mathcal{X}_{l}}\subseteq\mathbb{R}^{d_{l}}$, and $\sigma_{l}(W_{l}\cdots W_{2}\sigma_{1}(W_{1}\mathcal{X}_{0}))\subseteq\mathcal{X}_{l}\subseteq\mathbb{R}^{d_{l}}$ that satisfy $\mu_{\mathbb{R}^{d_{l}}}(\mathcal{X}_{l})>0$ and $\mu_{\mathbb{R}^{d_{l}}}(\tilde{\mathcal{X}}_{l})>0$.
Figure 2: Construction of $\mathcal{X}_{l}$ and $\tilde{X}_{l}$
Starting from $\mathcal{X}_{0}$, we recurrently construct $\tilde{\mathcal{X}}_{l}$ and $\mathcal{X}_{l}$ for $l=1,\ldots,L$. Since $W_{l}$ is injective, the space $W_{l}\mathcal{X}_{l-1}$ is $d_{l-1}$-dimensional. If $d_{l}>d_{l-1}$, then the measure $\mu_{\mathbb{R}^{d_{l}}}(W_{l}\mathcal{X}_{l-1})$ becomes 0, and setting $\tilde{\mathcal{X}_{l}}=W_{l}\mathcal{X}_{l-1}$ makes the analysis meaningless. Thus, we set a space $\mathcal{X}_{l}$ that includes $W_{l}\mathcal{X}_{l-1}$ and $\mu_{\mathbb{R}^{d_{l}}}(W_{l}\mathcal{X}_{l-1})>0$. Figure 2 schematically shows the construction of $\mathcal{X}_{l}$ and $\tilde{X}_{l}$. Let $\tilde{\mathcal{H}}_{l}=L^{2}(\tilde{\mathcal{X}}_{l})$, $\mathcal{H}_{l}=L^{2}(\mathcal{X}_{l})$, and $\eta_{l}(W_{l})=K_{W_{l}}$ be the Koopman operator from $\tilde{\mathcal{H}}_{l}$ to $\mathcal{H}_{l-1}$ with respect to $W_{l}$. In addition, let $A_{l}=K_{\sigma_{l}}$ be the Koopman operator from $\mathcal{H}_{l}$ to $\tilde{\mathcal{H}}_{l}$ with respect to an activation function $\sigma_{l}:\tilde{\mathcal{X}}_{l}\to\mathcal{X}_{l}$ that satisfies Assumption 1. Then, we have
@@
Let $D>0$ and $\mathcal{F}_{c}=\{F_{c}(\theta_{1},\ldots,\theta_{L},\cdot)\,\mid\,|\det{W_{1}^{*}W_{1}}|^{-1/4},\ldots,|\det{W_{L}^{*}W_{L}}|^{-1/4}\leq D\}$. Let
for $h\in\tilde{\mathcal{H}}_{l}$. This value depends on how large we set $\tilde{\mathcal{X}_{l}}$ compared with $W_{l}\mathcal{X}_{l-1}$, and by setting $\tilde{\mathcal{X}_{l}}$ sufficiently large, we can bound it by $1$ with a reasonable assumption (see Remark 17 for more details). In the same way as in Theorem 11, we obtain the following bound.
@@
Assume $p_{c,x}\in\mathcal{K}$ for $x\in\mathcal{X}_{0}$. Let $f_{l}=v\circ W_{L}\circ\sigma_{L-1}\circ\cdots\circ W_{l+1}\circ\sigma_{l}$. Then, we have
$\hat{R}(\mathcal{F}_{c},x_{1},\ldots,x_{S})\leq\sup_{|\det{W_{l}^{*}W_{l}}|^{-1/4}\leq D}\frac{{\color[rgb]{0,0,0}E(c)}\|v\|\prod_{l=1}^{L-1}\|A_{l}\|\alpha(f_{l})}{\sqrt{S}\prod_{l=1}^{L}|\det W_{l}^{*}W_{l}|^{1/4}},$ (4)